A market research company conducted a survey to find the percentage of people who exercise at least 30 minutes every day. Out of 186 persons surveyed, 57 replied that they exercise at least 30 minutes every day.

What is the standard error? Answer choices are rounded to the hundredths place.

  • a.)

    1.79

  • b.)

    0.21

  • c.)

    0.34

  • d.)

    0.03
Answer #1

\(S E=\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\)

Where,

\(\hat{p}=\frac{57}{186}=0.3065\)

Therefore,

\(S E=\sqrt{\frac{0.3065 * 0.6935}{186}}=0.03\)

d) \(0.03\)